OpenStudy (anonymous):

If y=3x-7, x>0, what is the minimum product of x^2y?

OpenStudy (michele_laino):

please, note that, your function is: \[f(x)=x ^{2}(3x-7)\] you have to explicit it, first step

OpenStudy (calculusfunctions):

This is an optimization problem. Let P represent the product. Then\[P =x ^{2}y\]If\[y =3x -7\]Then\[P =x ^{2}(3x -7)\]Now you should know how to use Calculus to solve the problem from here.

OpenStudy (michele_laino):

please use this theore: "A function f(x) is an increasing function if and only if the first derivative f'(x) is positive". I other words you have to find points x, at which your first derivative is positive

OpenStudy (michele_laino):

oops... theorem...

ganeshie8 (ganeshie8):

hmm x>0 can be a tricky constraint

OpenStudy (calculusfunctions):

 Is the graph of the product function, Thus you can see that the minimum product will be when 0 < x < 7/3.

OpenStudy (calculusfunctions):

Now use calculus to find the x (the critical number).

OpenStudy (michele_laino):

@calculusfunctions please note that the function of @chodefet is not definite for x<0

OpenStudy (calculusfunctions):

Yes I know that but thank you for your input anyway. I just drew the general graph. However if you you observe properly, you'll see I wrote that 0 < x < 7/3.

OpenStudy (michele_laino):

@calculusfunctions you are right!

OpenStudy (calculusfunctions):

Well @chodefet doesn't seem to be replying and so @Michele_Laino hopefully you can guide him without writing the solution for him because I will be logging out now.

OpenStudy (anonymous):

Let P=x²y, then P = x²y ..= x²(3x-7) ..= 3x³ - 7x²

OpenStudy (michele_laino):

@calculusfunctions ok! I try!

OpenStudy (anonymous):

Take the first derivative P'= 9x² - 14x ..= x(9x - 14) ..= 0 for x=0 and x=14/9

OpenStudy (anonymous):

okay so from here, do i just go through the process of derivative testing to find the minimum? thanks guys!

OpenStudy (anonymous):

Since x must be greater than zero, we are left to consider only x=14/9. The question now is: Does P have a minimum at x=14/9? Take the second derivative of P P''= 18x - 14

OpenStudy (anonymous):

thanks guys! i really appreciate the help!

OpenStudy (anonymous):

For x=14/9, P''= 18(14/9)-14 = 14 which is positive. Therefore P''>0 at x=14/9. So P is concave up at x=14/9 and thus P has a relative minimum at x=14/9

OpenStudy (anonymous):

Now, for x=14/9 y = 3x-7 y = 3(14/9)-7 y = -7/3 So, the minimum for x²y is P = x²y P = (14/9)²(-7/3) p = -5.646

OpenStudy (michele_laino):

@chodefet